OpenStudy (anonymous):

What is the side length of a cube with a volume of 0.064 m3? A. 0.000262 m B. 0.04 m C. 0.4 m D. 4 m @chelle,_clough

OpenStudy (anonymous):

to find the side length you can do V^.33 So... 0.064^.33=0.4

OpenStudy (anonymous):

Or instead of .33, 1/3 is equivalent

OpenStudy (anonymous):

let x be the side of cube. \[x^3=0.064=\frac{ 64 }{ 1000 }=\frac{ 4^3 }{ 10^3 }=\left( \frac{ 4 }{ 10 } \right)^3=\left( .4 \right) ^3\] x=?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

how do i fan u

OpenStudy (anonymous):

click on the man above left corner.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is above mathematics near 0